OpenStudy (anonymous):

Find an explicit rule for the nth term of the sequence. 1, 3, 9, 27, ...

OpenStudy (raden):

1 = 3^(1-1) 3 = 3^(2-1) 9 = 3^(3-1) 27 = 3^(4-1) ... ... so on what do u think about that ?

OpenStudy (anonymous):

so basically my equation would look like: \[a _{n} = 3 \times 1^{n-1}\] ?

OpenStudy (raden):



OpenStudy (anonymous):

Oh okay. Thank you!!

OpenStudy (raden):

you're welcome